OpenStudy (anonymous):

Solve 6 and two-ninths minus 4 and one-third. Show your work.

OpenStudy (anonymous):

write \[6\tfrac{2}{9}=\frac{9\times 6+2}{9}=\frac{56}{9}\]

OpenStudy (anonymous):

and \[4\tfrac{1}{3}=\frac{4\times 3+1}{3}=\frac{13}{3}\]

OpenStudy (anonymous):

satellite73 your answer is better

OpenStudy (anonymous):

then compute \[\frac{56}{9}-\frac{13}{3}=\frac{56}{9}-\frac{39}{9}\] subtract in the numerator

OpenStudy (anonymous):

mixed numbers are a pita

OpenStudy (anonymous):

can u type that out please

OpenStudy (anonymous):

if you are looking a bunch of code, refresh your browser and you should see the math